OpenStudy (anonymous):

derivitive of -4e^cosx

OpenStudy (amistre64):

-4 * e^cos(x) * -sin(x) 4sin(x) e^cos(x) is my best guess :)

OpenStudy (anonymous):

Wait is that e^(cosx*-sinx) or e^(cosx)*sinx???

OpenStudy (amistre64):

Dx(-4 e^cos(x)) ; -4 gets pulled out becuase it gets used in the end. For example; Dx(4x^2) = 4 Dx(x^2) = 4 (2x) = 8x. -4 Du(e^u) Dx(cos(x)) ; make u=cos(x) and use the chain rule. -4 (e^u) (-sin(x)) ; now we plug in the value of "u". -4 (e^cos(x)) (-sin(x)) ; now clean it up a little. 4sin(x) e^cos(x).